Find the Fourier Sine transform of f(x) = 1 in   0,l


1
Expert's answer
2020-12-25T15:22:43-0500

"\displaystyle
\mathcal{F}(f(x)) = \sqrt{\frac{2}{\pi}}\int_0^\infty f(x)\sin(\omega x) \, \mathrm{d}x\\

\begin{aligned}
\mathcal{F}(1) &= \sqrt{\frac{2}{\pi}}\int_0^l 1 \times \sin(\omega x) \, \mathrm{d}x = \sqrt{\frac{2}{\pi}}\int_0^l \sin(\omega x) \, \mathrm{d}x
\\&= \sqrt{\frac{2}{\pi}}\int_0^l \frac{-\cos(\omega x)}{\omega} \, \mathrm{d}x
\\&= \sqrt{\frac{2}{\pi}}\cdot \frac{-\cos(\omega x)}{\omega}\biggr\vert_0^l = \sqrt{\frac{2}{\pi}} \cdot \frac{\cos(\omega x)}{\omega}\biggr\vert_l^0
\\&= \sqrt{\frac{2}{\pi}}\left(\frac{1 - \cos(l\omega)}{\omega}\right)
\end{aligned}"